One type of sunburn occurs on exposure to UV light of wavelength in the vicinity of 305 nm. How many photons are in 1.15mJ burst of this radiation?
1
Expert's answer
2020-05-07T14:17:23-0400

"E=ch/\lambda" E - photon energy, с - speed of light (3*108m/s), "\lambda" - wavelength, h - Planck's constant (6,23*10-34kg*m2/s)

E=3*108m/s*6,23*10-34kg*m2/s/305*10-9m=6,12*10-19J

Whole E=0,00115J

N - photons ammount

N=0,00115J/6,12*10-19J=1,88*1015 photons

Answer:1,88*1015 photons